solve equation if x and y are whole numbers
xy=12

Answer by checkley71(8403)

HOW MANY SOLUTIONS DO YOU NEED?
X--Y
--------------
1*12=12
2*6=12
3*4=12
12*1=12
6*2=12
4*3=12
-1*-12=12
-2*-6=12
-3*-4=12
-12*-1=12
-6*-2=12
-4*-3=12